OpenStudy (anonymous):

nerdy girl in need!!!!! what are the solutions to 4(x-1)^2-16=0

OpenStudy (anonymous):

2(x-1) + 4 = 0 2(x-1) - 4 = 0

OpenStudy (lgbasallote):

what @GT did was difference of two squares..then equated the factors to 0

OpenStudy (experimentx):

4(x-1)^2-16 =0 =>4(x-1)^2 = 16 => (x-1)^2 = 16/4 => (x-1)^2 = 4 => 2^2 => x-1 = 2 OR x-1 = -2 => x = 2+1 or x = 1-2 => x=3 or x=-1

OpenStudy (callisto):

4(x-1)^2-16=0 Add 16 to both sides 4(x-1)^2-16 +16=0 +16 4(x-1)^2=16 Divide both sides by 4 4(x-1)^2 /4=16/4 (x-1)^2=4 take square root on both sides (x-1)= -2 or (x-1)= 2 x = -2+1 = -1 or x = 2+1 = 3

OpenStudy (anonymous):

x = 1 ± √ ( 16/4 ) = 3, -1